An event in the Summer Olympics is 10-meter springboard diving. In this event, the height s, in meters, of a diver above the water t seconds after jumping is given by s(t) = –4.9t2 + 7.1t + 10. What is the maximum height that the diver will be above the water? Round to the nearest tenth
Answer by Alan3354(69443)

An event in the Summer Olympics is 10-meter springboard diving. In this event, the height s, in meters, of a diver above the water t seconds after jumping is given by s(t) = –4.9t2 + 7.1t + 10. What is the maximum height that the diver will be above the water? Round to the nearest tenth
-----------------
s(t) = –4.9t2 + 7.1t + 10 is a parabola.
Max height is the vertex
The vertex is on the LOS (Line of Symmetry)
The LOS is t = -b/2a --> t = -7.1/-9.8 = 7.1/9.8
Max ht is s(7.1/9.8) meters